OpenStudy (anonymous):

Use the arc length formula and the given information to find r. Show your work for full credit. s = 8.1 ft θ = (3.14/3)rad r = ?

OpenStudy (jdoe0001):

\(\bf s = 8.1\qquad \qquad \theta = \cfrac{\pi}{3}\\ \quad \\ \quad \\ s = r\theta \implies \cfrac{s}{\theta} = r \) angle in radians

OpenStudy (anonymous):

I got 8.478

OpenStudy (jdoe0001):

hmm.... I got something else

OpenStudy (anonymous):

Ohh how'd you get it ?

OpenStudy (jdoe0001):

\(\bf \cfrac{8.1}{\frac{\pi}{3}} \implies \cfrac{8.1}{1} \times \cfrac{3}{\pi}\)

OpenStudy (anonymous):

76.302?

OpenStudy (jdoe0001):

\(\bf s = 8.1\qquad \qquad \theta = \cfrac{\pi}{3}\\ \quad \\ \quad \\ s = r\theta \implies \cfrac{s}{\theta} = r\\ \quad \\ \quad \\ \cfrac{8.1}{\frac{\pi}{3}} \implies \cfrac{8.1}{1} \times \cfrac{3}{\pi} \implies \cfrac{8.1 \times 3}{\pi}\) what would this give you?

OpenStudy (anonymous):

7.38

OpenStudy (jdoe0001):

hmm... I get 7.7388, but close enough I guess

OpenStudy (anonymous):

Thank you so much (:

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

Yeah I got that but my finger missed the 7 before 38 haha

OpenStudy (jdoe0001):

7.73493023426611331837 =)

OpenStudy (anonymous):

That works too (: